A chemist needs 12 liters of a 40% alcohol solution. She must mix a 20% solution and a 50% solution. How many liters of each well be required to obtain what she needs?
Answer by mananth(16946)

A chemist needs 12 liters of a 40% alcohol solution. She must mix a 20% solution and a 50% solution. How many liters of each well be required to obtain what she needs?
...
20% solution .............x liters
50% solution.............12-x liters
40%.......................12 liters (requirement)
...
I have not converted the percentage to 0.2,0.5,0.4.as generally done.
Even if we convert it, we again multiply by 100 for easy calculation
..
20*x+50*(12-x)=40*12
20x+600-50x=480
20x-50x=480-600
-30x=-120
/-30
x= 4 liters 20% solution
12-x = 8 liters of 50% solution.
...
m.ananth@hotmail.ca